please solve my problem: 12(3+y)=5(2y+8)
Answer by checkley77(12844)

12(3+y)=5(2y+8)
36+12y=10y+40
12y-10y=40-36
2y=4
y=4/2
y=2 ans.
Proof:
12(3+2)=5(2*2+8)
12*5=5(4+8)
60=5*12
60=60